# Eigenvector Localization and Universal Regime Transitions in Multiplex Networks: A Perturbative Approach

### v1

@@ VI Conclusions
In this work, we have deduced a first order perturbation approximation for the fictive coupling that characterizes the localised to delocalized transition for a social-contagion model on multiplex networks, $p^{*}$. With this approximation we also have obtained an analytical form for the non-dominant $IPR_{i}$ when $p<p^{*}$, coinciding with the numerical results in Ref. Ferraz de Arruda et al. (2020).
Furthermore, we have studied a dominance transition when varying $\gamma$ for one layer, mentioned before in Ref. Cozzo et al. (2013). We have given an analytical expression for the transition point, $\gamma^{*}$.
Our results completely characterize these regime transitions, giving a direct understanding of the mechanism causing it. However, the real nature of the layer-localized/delocalized transition is still elusive. In fact, structural transitions in multiplex networks are usually related to an eigenvalues crossing De Arruda et al. (2018); Gomez et al. (2013), while in this setting it seams not to be the case.
@@ Acknowledgements.
We acknowledge financial support from MINECO via Project No. PID2021-128005NB-C22 (MINECO/FEDER,UE), and Generalitat via Grant No. 2017SGR341

### v2

@@ VI Conclusions
In this work, we investigated the transition between layer-localized and delocalized regimes in a general contact-based contagion model on multiplex networks. Using a first-order perturbative analysis of the leading eigenvector of the supra-contact probability matrix, we derived an analytical expression for the fictive coupling $p^{*}$ that separates these regimes. This result not only reproduces the numerical scaling found in Ferraz de Arruda et al. (2020), but also provides a theoretical foundation for its universality, extending its validity beyond the SIS model to a broader class of contact-based dynamics.
We further obtained an analytical expression for the inverse participation ratio (IPR) of the non-dominant layer within the localized regime, confirming the power-law scaling with exponent $\alpha=4$ observed numerically. This scaling collapse confirms that the localized–delocalized crossover follows a universal functional form governed by the difference in layer spectral properties. This demonstrates that the observed universal behavior of the IPR has a spectral origin rooted in the perturbative structure of the supra-contact matrix.
In addition, we characterized the transition between dominant and non-dominant layers as a function of the intra-layer contact parameter $\gamma$. These results clarify the interplay between intra-layer activity and inter-layer dominance.
Finally, our numerical simulations confirmed the analytical results and revealed distinct dynamical signatures for the two transitions: the layer-localized/delocalized crossover and the dominance transition. The divergence patterns of the susceptibility peaks offer a clear phenomenological distinction between them.
Altogether, this work provides a unified spectral framework to understand localization and dominance transitions in multiplex contagion dynamics. While our perturbative approach captures the smooth nature of the observed crossover, a natural direction for future research is to explore whether, in suitable limits, the transition can become sharp, potentially associated with eigenvalue crossings or spectral gaps, typically associated with structural transitions in multiplex networks.
@@ Acknowledgements.
We acknowledge financial supportsupport from the Spanish grant PID2021-128005NB-C22, and from Generalitat de Catalunya underproject 2021-SGR-00856.